OpenStudy (anonymous):

Simplify 18 - 2[x + (x - 5)].

OpenStudy (anonymous):

18-2[2x-5] =18-4x+10 =-4x+28

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

no problem champ